OpenStudy (anonymous):

Integral of xe^(x/3)

OpenStudy (lgbasallote):

have you tried using integration by parts?

OpenStudy (anonymous):

Wouldn't u substitution work?

OpenStudy (anonymous):

did you try it? and succeed?

OpenStudy (anonymous):

U= x dV= e^(x/3) etc. try that and see if it works.

OpenStudy (anonymous):

Use integration by parts three times. ∫ x e^(x/3) dx let u = x : du = dx dv = e^(x/3) : v = 3e^(x/3) ∫ x e^(x/3) dx = 3x e^(x/3) - ∫ 3e^(x/3) dx = 3x e^(x/3) - 3 ∫e^(x/3) dx = 3x e^(x/3) - 9e^(x/3) + C